# Integrative Learning of Intensity Fluctuations of Quantum Dots under Excitation via a Tailored Mixture Hidden Markov Model

### v1

@@ Abstract
Semiconductor nano-crystals, known as quantum dots (QDs), have garnered significant interest in various scientific fields due to their unique fluorescence properties. One captivating characteristic of QDs is their ability to emit photons under continuous excitation. The intensity of photon emission fluctuates during the excitation, and such a fluctuation pattern can vary across different dots even under the same experimental conditions. What adding to the complication is that the processed intensity series are non-Gaussian and truncated due to necessary thresholding and normalization. As such, conventional approaches in the chemistry literature, typified by single-dot analysis of raw intensity data with Gaussian hidden Markov models (HMM), cannot meet the many analytical challenges and may fail to capture any novel yet rare fluctuation patterns among QDs. Collaborating with scientists in the chemistry field, we have developed an integrative learning approach to simultaneously analyzing intensity series of multiple QDs. Our approach still inherits the HMM as the skeleton to model the intensity fluctuations of each dot, and based on the data structure and the hypothesized collective behaviors of the QDs, our approach asserts that (i) under each hidden state, the normalized intensity follows a 0/1 inflated Beta distribution, (ii) the state distributions are shared across all the QDs, and (iii) the patterns of transitions can vary across QDs. These unique features allow for a precise characterization of the intensity fluctuation patterns and facilitate the clustering of the QDs. With experimental data collected on 128 QDs, our methods reveal several QD clusters characterized by unique transition patterns across three intensity states. The results provide deeper insight into QD behaviors and their design/application potentials.
Keywords: Clustering; Integrative learning; Multivariate time series; Zero-one inflation
@@ 1 Introduction
In the fields of chemistry and materials science, semiconductor Nano-crystals, commonly referred to as colloidal quantum dots (QDs), have gained significant attention in recent years. One captivating characteristic of the QDs is their ability to emit photons under photoexcitation, a phenomenon known as fluorescence. The applications of fluorescence property, particularly within the realm of optics and optoelectronics, have spurred extensive research and development efforts. For example, Li et al. (2018) developed a process to construct more efficient quantum LEDs and Bruns et al. (2017) used QDs as photostable probes to track and visualize biological tissues.
@@
The importance of the fluorescence property in applications stems from its abilities to elucidate the intrinsic properties and micro-environment of QDs, particularly at the single QD level. Beyond the aforementioned applications, a more comprehensive understanding of the fluorescence phenomenon would pave the way for designing QDs with specific purposes and providing essential guidelines for the synthesis of novel materials. In turn, this amplifies our focus on rigorous statistical analysis of fluorescence of single QDs, aiming to glean precise insights and optimize its implications.
@@
An intriguing phenomenon of fluorescence of single QDs is “Intensity Intermittency” or “Intensity Fluctuations”, which refers to the observation that the intensity of photon emission fluctuates during excitation, and such fluctuation patterns can vary across different QDs even under the same experimental conditions. Some patterns of intensity fluctuations (Nirmal et al., 1996) have been observed across various QD types and are well recognized such as Intensity Blinking, in which the intensity shows telegraph-like switching between high and low intensity states. Intensity Flickering has also been reported, in which the intensity exhibits frequent transitions across multiple intensity states. In reality, however, the intensity fluctuation patterns of single QDs are more complex and are not often described with scientific rigor. It is hypothesized that other patterns may exist. Notably, in our data, as depicted in Figure 1, the QDs in the first row are more likely to be of “Flickerin” style, the QDs in the second row exhibit more behaviors of “Blinking”, while the QDs in the last row show characteristics of neither fluctuation types.
@@
Figure 1: Intensity fluctuation patterns in quantum dots.
@@
In chemistry literature, researchers often conduct single-dot analysis of raw intensity series data, with, e.g., Gaussian hidden Markov models (HMM) (McKinney et al., 2006a). As the overall intensity level varies across QDs, the resulting high, medium, and low intensity states from HMMs could also vary across QDs. Their selection, labeling, and correspondence can then be arbitrary and often rely on the experiences of the researcher. In more recent studies, some advanced analytical methods have been developed, encompassing probability distribution analysis (Efros and Rosen, 1997; Kuno et al., 2000), burst variance analysis (Frantsuzov et al., 2008), FRET two-kernel density estimator (Sisamakis et al., 2010), and fluorescence correlation spectroscopy (Magde et al., 1972; Mücksch et al., 2018). While each of these methods focuses on elucidating the fluctuation patterns of individual QDs, they tend to overlook the aspect of similarity across different QDs. As such, they can’t meet the many analytical challenges and may fail to capture any novel yet rare fluctuation patterns among QDs.
@@
Figure 2: Proposed pipeline of quantum dot analysis from data collection, data processing, to integrative analysis with MHMM.
@@
Collaborating with scientists in chemistry, we innovate an integrative learning pipeline (as shown in Figure 2) to simultaneously analyze standardized and robustified intensity series of multiple QDs. To achieve this, we first develop a data pre-processing procedure based on the scientific experiment, by eliminating background noise from the intensity values and thresholding and normalizing raw intensity values to remove the inherent variation between different QDs. After rescaling, the intensity values in each series then become bounded between 0 and 1 and can be exactly either 0 or 1. At the end of the pipeline, the processed intensities are analyzed integratively using our proposed approach, where the state identification and clustering are performed iteratively.
@@
Our analytic approach still inherits the HMM as the skeleton, that is, for each dot, we assume the intensity fluctuations are governed by a set of unobserved hidden states that operate as a Markov chain. We further assume that (1) under each state, the standardized intensity follows a 0/1 inflated Beta distribution based on empirical evidence, (2) the hidden state distributions are shared across all the QDs to enable integrative learning, and (3) the patterns of transitions can vary across QDs thus giving rise to a mixture HMM (MHMM). These features lead to a precise characterization of the intensity fluctuation patterns and facilitate an objective clustering of the QDs.
@@
The remainder of the paper is organized as follows. In Section 2, we describe the data and the problem setup for studying the intensity fluctuations. In Section 3, after an overview of existing methods, we propose the MHMM model with inflated Beta distribution and briefly discuss its computation and theoretical properties. Extensive simulation studies are presented in Section 4. In Section 5, we thoroughly discuss the results of analyzing our QD dataset. In Section 6, we provide a few concluding remarks and discuss future works.
2 Data and Pre-Processing Pipeline
@@
The data were compiled from 128 quantum dot samples. For each QD, we obtained 2000 intensity measurements across a 100-second time frame. In the following, we detail the data acquisition and processing methodology.
@@ 2.2 Pre-Processing
    Thresholding of extreme values: For each QD, the intensity threshold is set as the average of its 90th percentile and its maximum background-free intensity. Intensity values below zero are reset to zero, while those above the threshold were capped at the threshold value.
@@
We stress that the data preprocessing procedure described above is based on the nature of the chemical experiments and is deemed justifiable by our close collaborators in chemistry. This procedure yields a standardized dataset where each intensity series consists of 2000 equally-spaced intensity values for a 100-second time frame and with all the values ranging from 0 to 1 (inclusive). This makes a direct comparison across the QDs possible.
Table 1 presents some summary statistics for standardized intensity measurements. It indicates that the mean intensity values of the QDs are relatively concentrated, with a median of 0.694 and an interquartile range of 0.614 to 0.774. Additionally, the zero rate and the one rate remain low, with a maximum of 0.021, suggesting very limited occurrences of extreme values in the intensity values. These confirm that the standardized intensity values are comparable across the QDs. Furthermore, recall that Figure 1 shows that QDs may exhibit distinctive behaviors of intensity fluctuation over time. We therefore aim to perform an integrative learning of all QDs to reveal the unique yet interrelated intensity fluctuation mechanisms and potential clusters of the QDs.
Table 1: Summary statistics of standardized intensity series.
@@
3 Integrative Learning of Intensity Fluctuations of QDs
@@ 3.1 Overview
In the chemistry literature, researchers often analyze the dynamics and intensity fluctuations of each single QD. For example, McKinney et al. (2006b) introduced an HMM approach, assuming that different trajectories shared a homogeneous transition pattern. Consequently, their approach involved fitting an HMM model to each (representative) trajectory and calculating the average transition matrices. Pirchi et al. (2016) utilized an $\mbox{H}^{2}\mbox{MM}$ framework, an enhanced version of HMM. The novelty of their approach lied in the consideration of hidden state transitions and non-transition times; two sets of hidden states were assigned – one for transitions and another for non-transitions. These HMM based single-dot approaches allow for the classification of intensity levels into discrete states, such as high, medium, and low. However, challenges arise because intensity levels vary across QDs, leading to subjective labeling and reliance on researcher expertise for state correspondence. While HMMs provide valuable insights into individual QDs, they lack the capacity to analyze patterns across multiple QDs simultaneously, missing opportunities to identify shared behaviors or rare but novel anomalies.
To address data heterogeneity, finite mixture models (Dempster et al., 1977) have been widely adopted in various fields, including medicine (Schlattmann and Böhning, 1993), public health (Fahey et al., 2011), psychology (Steinley and Brusco, 2011), and economics (Deb et al., 2011; Deb and Trivedi, 1997). These models decompose data into a weighted sum of distributions, each representing a cluster, thereby capturing group-level patterns. The integration of HMMs and mixture models, particularly in the form of Mixture Hidden Markov Models (MHMMs), has shown promise in addressing temporal and cluster-level complexities. MHMMs extend HMMs by allowing the transition dynamics to vary across clusters, enabling simultaneous state identification and clustering. Historically, MHMMs have been mainly applied to Gaussian data, such as in neuroscience to classify EEG signals (Wang et al., 2018) or in economics to study market volatility (Dias et al., 2010). However, Gaussian assumptions are often unsuitable for datasets like QD intensities, where distributions exhibit bounded behavior and inflation at extreme values (e.g., 0 and 1).
3.2 Mixture HMM with Inflated Beta
@@
Our approach inherits the HMM as the skeleton, that is, for each dot, we assume the intensity fluctuations are governed by a set of unobserved hidden states that operate as a Markov chain. With $M$ states, the state of the intensity $x_{i,t}$ can be represented as a vector $\boldsymbol{s}_{i,t}=(s_{i,t,1},\ldots,s_{i,t,M})^{\rm T}$, where $s_{i,t,h}$, $h=1,\ldots,M$, are indicator variables and $s_{i,t,h}=1$ only if $x_{i,t}$ is at state $h$. We use $\boldsymbol{s}_{i}=(\boldsymbol{s}_{i,1},\ldots,\boldsymbol{s}_{i,T})^{\rm T}$ to represent the state series of $\boldsymbol{x}_{i}$ and use $\boldsymbol{S}=(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{N})^{\rm T}$ to collect state series for all the intensity series.
@@
It is now in order to capture the heterogeneity in the intensity fluctuation across the QDs. We achieve this by considering the transition patterns in the hidden Markov process and assuming that the state transition probabilities can vary across QDs. This gives rise to a mixture HMM (MHMM), with which all QDs can be clustered into $K$ clusters/groups, each characterized by a unique state transition or fluctuation pattern. Specifically, let the cluster number of each dot $i$ be represented by a vector $\boldsymbol{z}_{i}=(z_{i,1},\ldots,z_{i,K})^{\rm T}$, where $z_{i,k}=1$ if dot $i$ belongs to group $k$ and otherwise $z_{i,k}=0$, for $k=1,\ldots,K$, $i\in\{1,\ldots,N\}$, and let $\boldsymbol{Z}=(\boldsymbol{z}_{1},\ldots,\boldsymbol{z}_{N})^{\rm T}$. Let $\delta_{k}=Pr(z_{i,k}=1)$, $k=1,\ldots,K$, with $\sum_{k=1}^{K}\delta_{k}=1$. For each cluster, we use $\boldsymbol{\Pi}_{k}$ to denote the transition matrix, where its element $(\boldsymbol{\Pi}_{k})_{p,q}$ gives the probability of switching to state $q$ from state $p$, i.e., $(\boldsymbol{\Pi}_{k})_{p,q}=Pr(s_{i,t,q}=1\mid s_{i,t-1,p}=1)$. We also use $\boldsymbol{\Pi}$ to denote the collection of all the transition matrices $\boldsymbol{\Pi}_{k},k\in\{1,\ldots,K\}$.
We term the proposed method as the integrative mixture hidden Markov model with 0/1 inflated Beta, denoted as MHMM-$\beta$. The tailored model structures imposed above lead to a precise characterization of the intensity fluctuation patterns and facilitate an objective clustering of the QDs. To summarize and visualize, Figure 3 shows a conceptual diagram of the proposed MHMM-$\beta$ model. The clusters share the same set of three hidden states and are distinguished by three different transition patterns. The three shared states, i.e., the three inflated Beta distributions, are shown by the three histograms at the bottom of the figure. The potential transitions are represented by the arrows with the thickness indicating the magnitude of the probability. A QD in Cluster 1 tends to stay in its current state and rarely transits to other states, while a QD in Cluster 2 may always transit to the high state from the low state (essentially it means that the QD does not spend any time in the low state).

### v2

# Integrative Learning of Quantum Dot Intensity Fluctuations under Excitation via Tailored Dynamic Mixture Modeling

@@ Abstract
Semiconductor nano-crystals, known as quantum dots (QDs), have attracted significant attention for their unique fluorescence properties. Under continuous excitation, QDs emit photons with intricate intensity fluctuation: the intensity of photon emission fluctuates during the excitation, and such a fluctuation pattern can vary across different QDs even under the same experimental conditions. What adding to the complication is that the processed intensity series are non-Gaussian and truncated due to necessary thresholding and normalization. Conventional normality-based single-dot analysis fall short of addressing these complexities. In collaboration with chemists, we develop an integrative learning approach to simultaneously analyzing intensity series from multiple QDs. Motivated by the unique data structure and the hypothesized behaviors of the QDs, our approach leverages the celebrated hidden Markov model as its structural backbone to characterize individual dot intensity fluctuations, while assuming that, in each state the normalized intensity follows a 0/1 inflated Beta distribution, the state/emission distributions are shared across the QDs, and the state transition dynamics can vary among a few QD clusters. This framework allows for a precise, collective characterization of intensity fluctuation patterns and have the potential to transform current practice in chemistry. Applying our method to experimental data from 128 QDs, we reveal three shared intensity states and capture several distinct intensity transition patterns, underscoring the effectiveness of our approach in providing deeper insights into QD behaviors and their design and application potential.
Keywords: Clustering; EM algorithm; Hidden Markov model; Integrative analysis; Zero-one inflation
@@ 1 Introduction
In the fields of chemistry and material science, semiconductor Nano-crystals, commonly referred to as colloidal quantum dots (QDs), have gained significant attention in recent years. One captivating characteristic of the QDs is their ability to emit photons under photoexcitation, a phenomenon known as fluorescence. The applications of fluorescence property, particularly within the realm of optics and optoelectronics, have spurred extensive research and development efforts. For example, Li et al. (2018) developed a process to construct more efficient quantum LEDs, and Bruns et al. (2017) used QDs as photostable probes to track and visualize biological tissues.
@@
The importance of the fluorescence property in applications stems from its abilities to elucidate the intrinsic properties and micro-environment of QDs, particularly at the single QD level. Beyond the aforementioned applications, a more comprehensive understanding of the fluorescence phenomenon would pave the way for designing QDs with specific purposes and providing essential guidelines for the synthesis of novel materials. This urgent need amplifies our focus on a comprehensive and rigorous statistical analysis of the fluorescence of the QDs.
@@
An intriguing phenomenon of the fluorescence of QD is the so-called “Intensity Intermittency” or “Intensity Fluctuation”, which refers to the observation that the intensity of photon emission fluctuates during the excitation process, and such fluctuation patterns can vary across different QDs even under the same experimental conditions. Some dynamics or patterns of intensity fluctuations (Nirmal et al., 1996) have been observed across various QD types and have been well recognized in chemistry literature. For example, “Intensity Blinking” describes the pattern in which the intensity shows telegraph-like switching between high and low intensity states. “Intensity Flickering” has also been reported, in which the intensity exhibits frequent transitions across multiple intensity states. In reality, however, the intensity fluctuation patterns of single QDs are more complex and are not often described with scientific rigor. It is hypothesized that other patterns may exist. Notably, in our data, as depicted in Figure 1, while some QDs exhibit clear behaviors of either “Blinking” or “Flickering”, others may show characteristics of neither of these two known fluctuation types.
@@
Figure 1: Quantum dot data: Intensity fluctuation patterns in from some selected quantum dots.
@@
In chemistry literature, researchers often conduct single-dot analysis of raw intensity series data, with, e.g., Gaussian hidden Markov models (HMM) (McKinney et al., 2006). As the overall intensity level varies across QDs, the resulting high, medium, and low intensity states from HMMs could also vary across QDs. Their selection, labeling, and correspondence can then be arbitrary and often rely on the experiences of the researcher.
In more recent studies, some advanced analytical methods have been developed, encompassing probability distribution analysis (Efros and Rosen, 1997; Kuno et al., 2000), burst variance analysis (Frantsuzov et al., 2008), FRET two-kernel density estimator (Sisamakis et al., 2010), and fluorescence correlation spectroscopy (Magde et al., 1972; Mücksch et al., 2018). While each of these methods focuses on elucidating the fluctuation patterns of individual QDs, they tend to overlook the aspect of similarity across different QDs. As such, they fall short in meeting the many analytical challenges and may fail to capture any novel yet rare fluctuation patterns among QDs.
@@
Figure 2: Proposed pipeline of quantum dot analysis from data collection, data processing, to integrative learning with dynamic mixture model.
@@
Closely collaborating with scientists in chemistry, we innovate an integrative learning pipeline (as shown in Figure 2) to simultaneously analyze standardized and robustified intensity series of multiple QDs. To achieve this, we first develop a data pre-processing procedure based on the nature of the experiment and limitations of the equipment, by eliminating background noise from the intensity values and thresholding and normalizing raw intensity values to remove the inherent variations between different QDs. After rescaling, the intensity values in each series then become bounded between 0 and 1 and can also be exactly equal to either 0 or 1. These processed intensities are then analyzed using our proposed integrative learning approach, in which state identification, transition pattern recognition, and QD clustering are performed simultaneously.
@@
Our analytic approach still inherits the celebrated HMM as the skeleton, that is, for each dot, we assume the intensity fluctuations are governed by a set of unobserved hidden states that operate as a Markov chain. To enable integrative learning, we further assume that (1) under each state, the standardized intensity follows a 0/1 inflated Beta distribution based on empirical evidence, (2) the hidden state distributions are shared among all the QDs, and (3) the patterns of transitions can vary across QDs thus giving rise to a mixture HMM (MHMM). These features lead to a precise and collective characterization of the intensity fluctuation patterns and facilitate an objective clustering of the QDs.
@@
The remainder of the paper is organized as follows. In Section 2, we describe the data and the problem setup for studying the intensity fluctuations. In Section 3, after an overview of existing methods, we propose a dynamic mixture model with inflated Beta distribution, develop an efficient computational algorithm for parameter estimation, and discuss issues related to practical implementation, model selection, and inference. Extensive simulation studies are presented in Section 4. In Section 5, we thoroughly analyze the QD dataset and discuss the results and implications. In Section 6, we provide a few concluding remarks and discuss future research directions.
2 Data Description and Pre-Processing Pipeline
@@
The data were compiled from 128 quantum dot samples. For each QD, we obtained 2000 equally-spaced intensity measurements of photon emission over a 100-second time frame (under continuous excitation). In the following, we detail the data acquisition and processing methodology.
@@ 2.2 Pre-Processing
    Thresholding of extreme values: For each QD, the intensity threshold is set as the average of its 90th percentile and its maximum background-free intensity. Intensity values below zero are reset to zero, while those above the threshold are capped at the threshold value.
@@
We stress that the data preprocessing procedure described above is based on the nature of the chemical experiments and the limitations of the experimental equipment, and it is justified from the chemistry perspectives. This procedure yields a standardized dataset where each intensity series consists of 2000 equally-spaced intensity values for a 100-second time frame and with all the values ranging from 0 to 1 (inclusive). This makes a direct comparison across the QDs meaningful and thus enables their integrative learning.
Table 1 presents some summary statistics for the standardized intensity measurements. The results indicate that the mean intensity values of the QDs (averaged over time for each QD) are relatively concentrated, with a median of 0.694 and an interquartile range of 0.613 to 0.774. Additionally, the zero rate and the one rate remain low, with a maximum of 0.038, suggesting very limited occurrences of extreme values in the intensity values. These confirm that the standardized intensity values are comparable across the QDs. Furthermore, recall that Figure 1 shows that QDs may exhibit distinctive behaviors of intensity fluctuation over time. We therefore aim to perform an integrative learning of all QDs to reveal the unique yet interrelated intensity fluctuation mechanisms and potential clusters of the QDs.
Table 1: Quantum dot data: Summary statistics of standardized intensity series.
@@
3 Integrative Dynamic Mixture Model of Intensity Fluctuations
@@ 3.1 Overview
In the chemistry literature, researchers often analyze the dynamics and intensity fluctuations of each single QD. For example, McKinney et al. (2006) introduced an HMM approach, assuming that different trajectories shared a homogeneous transition pattern. Consequently, their approach involved fitting an HMM model to each (representative) trajectory and calculating the average transition matrices. Pirchi et al. (2016) utilized an $\mbox{H}^{2}\mbox{MM}$ framework, an enhanced version of HMM. The novelty of their approach lied in the consideration of hidden state transitions and non-transition times; two sets of hidden states were assigned – one for transitions and another for non-transitions.
These HMM based single-dot approaches allow for the classification of intensity levels into discrete states, such as high, medium, and low. However, challenges arise because intensity levels vary across QDs, leading to subjective labeling and reliance on researcher expertise for state correspondence. While HMMs provide valuable insights into individual QDs, they lack the capacity to analyze patterns across multiple QDs simultaneously, missing opportunities to identify shared behaviors or rare but novel anomalies.
To address data heterogeneity, finite mixture models (Dempster et al., 1977; Yao and Xiang, 2024) have been widely adopted in various fields, including medicine (Schlattmann and Böhning, 1993), public health (Fahey et al., 2011), genetics (Chen et al., 2018), psychology (Steinley and Brusco, 2011), and economics (Deb et al., 2011; Deb and Trivedi, 1997). These models decompose data into a weighted sum of distributions, each representing a cluster, thereby capturing group-level patterns.
The integration of HMMs and mixture models, particularly in the form of Mixture Hidden Markov Models (MHMMs), has shown promise in addressing temporal and cluster-level complexities. MHMMs extend HMMs by allowing the transition dynamics to vary across clusters, enabling simultaneous state identification and clustering. Historically, MHMMs have been mainly applied to Gaussian data, such as in neuroscience to classify EEG signals (Wang et al., 2018) or in economics to study market volatility (Dias et al., 2010). However, Gaussian assumptions are often unsuitable for datasets like QD intensities, where distributions exhibit bounded behavior and inflation at extreme values (e.g., 0 and 1).
3.2 Mixture HMM with 0/1 Inflated Beta
@@
Our approach inherits the HMM as the skeleton, that is, for each dot, we assume the intensity fluctuations are governed by a set of unobserved hidden states that operate as a Markov chain. With $M$ states, the state of the intensity $x_{i,t}$ can be represented as a vector $\boldsymbol{s}_{i,t}=(s_{i,t,1},\ldots,s_{i,t,M})^{\rm T}$, where $s_{i,t,h}$, $h=1,\ldots,M$, are indicator variables and $s_{i,t,h}=1$ only if $x_{i,t}$ is at state $h$.
We use $\boldsymbol{s}_{i}=(\boldsymbol{s}_{i,1},\ldots,\boldsymbol{s}_{i,T})^{\rm T}$ to represent the state series of $\boldsymbol{x}_{i}$ and use $\boldsymbol{S}=(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{N})^{\rm T}$ to collect state series for all the intensity series.
@@
It is now in order to capture the heterogeneity in the intensity fluctuation across the QDs. We achieve this by considering the transition patterns in the hidden Markov process and assuming that the state transition probabilities can vary across QDs. This gives rise to a mixture HMM (MHMM), with which all QDs can be clustered into $K$ clusters/groups, each characterized by a unique state transition or fluctuation pattern. Specifically, let the cluster number of each dot $i$ be represented by a vector $\boldsymbol{z}_{i}=(z_{i,1},\ldots,z_{i,K})^{\rm T}$, where $z_{i,k}=1$ if dot $i$ belongs to group $k$ and otherwise $z_{i,k}=0$, for $k=1,\ldots,K$, $i\in\{1,\ldots,N\}$, and let $\boldsymbol{Z}=(\boldsymbol{z}_{1},\ldots,\boldsymbol{z}_{N})^{\rm T}$. Let $\delta_{k}=P(z_{i,k}=1)$, $k=1,\ldots,K$, with $\sum_{k=1}^{K}\delta_{k}=1$. For each cluster, we use $\boldsymbol{\Pi}_{k}$ to denote the transition matrix, where its element $(\boldsymbol{\Pi}_{k})_{p,q}$ gives the probability of switching to state $q$ from state $p$, i.e., $(\boldsymbol{\Pi}_{k})_{p,q}=P(s_{i,t,q}=1\mid s_{i,t-1,p}=1)$. We then use $\boldsymbol{\Pi}$ to denote the collection of all the transition matrices $\boldsymbol{\Pi}_{k},k\in\{1,\ldots,K\}$.
We term the proposed method as the integrative mixture hidden Markov model with 0/1 inflated Beta, denoted as MHMM-$\beta$. The tailored model structures imposed above lead to a precise and collective characterization of the intensity fluctuation patterns and facilitate an objective clustering of the QDs. To summarize and visualize, Figure 3 shows a conceptual diagram of the proposed MHMM-$\beta$ model. The clusters share the same set of hidden states ($M=3$) and are distinguished by different transition patterns ($K=3$). The shared states, i.e., the three inflated Beta distributions, are shown by the three histograms at the bottom of the figure. The potential transitions are represented by the arrows, where the thickness indicates the magnitude of the corresponding probability. A QD in Cluster 1 tends to stay in its current state and rarely transits to other states, whereas a QD in Cluster 2 may always transit to the high state from the low state (essentially it means that the QD does not spend any time in the low state).